OpenStudy (anonymous):

Would x=1? 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

And what is y?

OpenStudy (phi):

yes. add the 2 equations. you get 2x -15x + 5y -5y = -3 + -10 -13 x = -13 x = -13/-13 = 1 btw, y = -1

OpenStudy (anonymous):

Thanks!

OpenStudy (phi):

2x + 5y = -3 plop in 1 for x 2*1 +5y = -3 2+5y= -3 2+ -2 + 5y = -3 + -2 5y = -5 y = -5/5 =-1